OpenStudy (anonymous):

Calculate the lim h-->0 sin9h/h.

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{\lim_{x\to0}\frac{\sin x}{x}=1}\]Remember this identity?

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{h\to0}\frac{\sin9h}{h}\cdot \frac{9}{9}\]Multiply top and bottom by 9,

zepdrix (zepdrix):

\[\Large\bf\sf 9\lim_{h\to0}\frac{\sin(9h)}{(9h)}\]

zepdrix (zepdrix):

See how we can apply the identity from here? :)

OpenStudy (anonymous):

try to multply and divide by 9,then do 9h =u and take the limit when u goes to zero

OpenStudy (anonymous):

Ohh okay! I was looking at a solution manual and I couldn't figure out why they had a 9h at the denominator but this definitely helped, thank you. :)

zepdrix (zepdrix):

yay team \c:/